OpenStudy (anonymous):

Solve the following system of equations. x + 4y + z = –10 3x – 3y + 6z = –21 x + 2y + 2z = –10 Answer (–4, –1, –2) (–3, –2, –3) (1, –2, –3) (2, –2, –4)

OpenStudy (anonymous):

There are many ways of solving this. a, Substitution b, Elimination c, Using a matrix. Which one do you prefer, I can show it for you.

OpenStudy (anonymous):

a

OpenStudy (anonymous):

For that you have to express one of the equations in the form of x=somthing(no x in it) Than you have to put this back to the other two equations.

OpenStudy (anonymous):

from the first one we can see that x=-10-4y-z Is it clear until this?

OpenStudy (anonymous):

The second and third equations are: 3x – 3y + 6z = –21 x + 2y + 2z = –10 Now instead of x you write -10-4y-z So it will be 3(-10-4y-z)-3y+6z=-21 -10-4y-z+2y+2z=-10 Clear?

OpenStudy (anonymous):

my preferred method is cheating, but it will not help you on an exam

OpenStudy (anonymous):

cheating? :) How?

OpenStudy (anonymous):

got it with copy and paste and commas between the equations.